OpenStudy (anonymous):

What is the slope of the line 2x - 5y = -10?

OpenStudy (anonymous):

Convert it into a form y = mx + b. m will be the slope.

OpenStudy (anonymous):

2x - 5y = -10 => -5y = -2x - 10 => 5y = 2x + 10 => y = 2/5x + 2 So, slope is 2/5.

OpenStudy (anonymous):

woow hes smart

OpenStudy (anonymous):

Hey GT can you answer one of the Questions i posted please